OpenStudy (chris215):

If f is a function such that lim x->a f(x)-f(a)/x-a =7 , then which of the following statements must be true? a. f(a) = 7 b. f(x) is continuous at x = 7 c. the linear approximation for f(x) at x = a is 7 d. f '(a) = 7

OpenStudy (anonymous):

i think its the derivative of a

OpenStudy (chris215):

is that d?

OpenStudy (freckles):

is that suppose to say \[\lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a}=7 ?\]

OpenStudy (chris215):

yes

OpenStudy (freckles):

Here are a few hints: 1) recall definition of derivative 2) if derivative exist at x=a, then the function is continuous at x=a 3) the linear approximation to f(x) at x=a is given by L(x)=f'(a)(x-a)+f(a) and what happens if you use the linear approximation to evaluate L(a)

OpenStudy (chris215):

the linear approximation for f(x) at x = a is 7

OpenStudy (freckles):

L(x)=f'(a)(x-a)+f(a) Finding L(a) <--we are finding the linear approximation at x=a L(a)=f'(a)(a-a)+f(a) L(a)=?

OpenStudy (freckles):

If you are saying "the linear approximation for f(x) at x = a is 7" is your answer, I'm kind asking you how do you know that f(a)=7 since the linear approximation to f at x=a is L(a)=f(a) we do not know f(a)=7 so we cannot say this choice

OpenStudy (freckles):

So use the other two hints given

OpenStudy (freckles):

what is your definition of derivative?

OpenStudy (freckles):

you are almost done weeding out the false statements

OpenStudy (freckles):

I must go good luck